Jims student loans totaled $10,000.Part was a plus loan made at 3% interestand the rest was a sally mae made at 5% interest. After one year jims loans accumlated $380 in interest. What was the original amount of each loan ?
Answer by math_helper(2461)

Let p = amount of plus loan
and s = amount of sally mae loan
p + s = 10000         (1)
0.03p + 0.05s = 380   (2)

Let's eliminate p by careful choice of multiplication of (1).
Multiply (1) by 0.03:
0.03p + 0.03s = 300

Subtract (2) from (1)
  0.03p + 0.03s = 300
-0.03p - 0.05s  = -380
--——————————————————————
         -0.02s = -80
  s = -80/-0.02 = 4000

From (1)  p + 4000 = 10000  so p = 6000

—

Ans:
          plus loan was $6000
          sallie mae loan was $4000
—

Check:   6000(0.03) + 4000(0.05) = 
              180 + 200  = 380  (ok, interest matches)